The charges q, 3.0 10 C -6.5 10C, and1.910Careed at the cornersof +y-axis is up along the page.) 5.0 m 3.0 m 4.0 m 93
Answer #1

Force on q1 due to charge q2 which will be given as -

F21 = ke |q2| |q1| / r212

F21 = [(9 x 109 Nm2/C2) (6.5 x 10-7 C) (3.8 x 10-7 C)] / (5 m)2

F21 = 8.89 x 10-5 N

Force on q1 due to charge q3 which will be given as -

F31 = ke |q3| |q1| / r312

F31 = [(9 x 109 Nm2/C2) (1.9 x 10-7 C) (3.8 x 10-7 C)] / (3 m)2

F31 = 7.22 x 10-5 N

The x-component of force on q1 which will be given as -

F1,x = F21,x + F31,x

F1,x = (8.89 x 10-5 N) cos 600 + (7.22 x 10-5 N) cos 900

F1,x = (8.89 x 10-5 N) (0.5)

F1,x = 4.445 x 10-5 N

The y-component of force on q1 which will be given as -

F1,y = F21,y + F31,y

F1,y = (8.89 x 10-5 N) sin 600 + (7.22 x 10-5 N) sin 900

F1,y = [(8.89 x 10-5 N) (0.866) + (7.22 x 10-5 N)]

F1,y = 1.491 x 10-4 N

Therefore, the force on charge q1 which will be given as -

Fon 1 = \sqrt{}F1,x2 + F1,y2

Fon 1 = \sqrt{}(4.445 x 10-5 N)2 + (1.491 x 10-4 N)2

Fon 1 = \sqrt{}2.420 x 10-8 N2

Fon 1 = 1.55 x 10-4 N